A package contains 12 resistors, 3 of which are defective. If 3 are selected, find the probability of getting

(a) no defective resistor.

(b) one defective resistor.

(c) all defective resistors 


1
Expert's answer
2022-04-11T18:08:04-0400

The probability for a defective resistor is,P(D)=3/12=0.25

The probability for a non-defective resistor is 2-0.25=0.75

Find the probabilities for defective resistors from 0 to 3 out of 3 total

"P(0)=1(0.75)^3(0.25)^0=0.422"

"P(1)=3(0.75)^2(0.25)^1=0.422"

"P(2)=3(0.75)^1(0.25)^2=0.141"

"P(3)=1(0.75)^0(0.25)^3=0.016"

a. P(0)=0.422

b. P(1)=0.422

c.P(3)=0.016